# Pz Cats: Photometric redshift catalogs based on DES Y3 BAO sample

### v1

@@ 4.2 Precision with respect to the redshift
Complementary to the results above, we show $\sigma_{68}$ w.r.t. the spectroscopic redshift of the matched galaxies used for testing (all galaxies for BPZ). The result is shown in Figure 10. BPZ in figure 10(a) looses precision when $z>1.0$, for ENF/DNF(figs. 10(c),10(d)) this happens at $z>0.9$. The only photo-z estimator to show consistent and precise results is ANNz2.
@@
Figure 10: $\sigma^{68}$ for each $z_{spec}$ bin. Blue: positive error. Red: negative error.
@@
Table 4: Catastrophic results for each algorithm.
We compared the relative frequency of galaxies classified with Gaussian PDFs from the testing sample for each algorithm. In Figure 11, we show the number of Gaussians ($N_{Gauss}$) up to $|\sigma_{z}|$ compared to the number of galaxies in the test sample up to $|\sigma_{z}|$. BPZ has nearly 70% Gaussian PDFs with bias bigger than 0.075 and ANNz2 has nearly 60%. ENF reaches a maximum of 25% of Gaussian PDFs for very small bias, after $|\sigma_{z}|>0.025$, the frequency saturates and only 15% of the PDFs are nearly Gaussian. The opposite happens to DNF, it is an improvement compared to ENF, but saturation is below 25%. This result shows an important situation that may occur, even though the PDFs are smooth, there are catastrophic redshift results.
Figure 11: Relative frequency of Gaussian distributions compared to the total number of testing samples w.r.t. the absolute bias ($\sigma_{z}$) for each estimator.
@@ 4.3 Filters with respect to the bias
Figure 12: Bias distribution for each filter. Red: DNF, black: BPZ, gray: ENF and blue: ANNz2.
@@ 4.4 Colour cut
In the previous test, we based our $\sigma_{68}$ analysis w.r.t. the colour coverage of our spectroscopic sample, in Figure 1. For each estimator, we cut the redshifts restriction to the colours $41<g-i<$ 4 and $1.40.5<r-i<$ 1.4. Then, we selected the redshifts whose $\sigma_{68}$ is smaller than $10\%$ of the photo-z value. In table 5, there are the $z_{eff}$ considering the colour cut and it is clear that they are compatible with the entire sample photo-z estimation. lastly, the number of objects reduces significantly to DNF, so depending on the use of the catalog, one must understand that an accurate representation of the filters by the training sample may reduce statistics.
Estimator $z_{eff}$ number of galaxies
@@
We checked the bias distribution normalized by $\sigma_{68}$ in terms of the faintest part of each filter. This is shown in Figure 12, the only algorithm that is symmetric over the bias is ANNz2, BPZ underestimates the faint galaxies, while DNF and ENF overestimate them.
@@ 5 Summary
We selected 25,760 galaxies from four different spectroscopic surveys and cross-matched them with the DES Y3 BAO sample. These galaxies served to understand the redshift bias $\sigma$ and its 68th percentile $\sigma_{68}$ and are representative in terms of magnitude. We found that within a range of $0.850.79<z_{p}<$ 0.85 there is the lowest $\sigma$ for all the estimators we analysed. DNF has the biggest absolute value of the bias, while ENF, ANNz2 and BPZ lose precision for a redshift range below 0.7 and higher than 0.9.
@@
We also investigated the $\sigma_{68}$ metric, a robust metric to estimate the error of the algorithms. ANNz2 and BPZ showed consistent $\sigma_{68}$ for all bins, while ENF and DNF have an asymmetric error distribution, with lower negative error values than the positive ones and a considerable error increase for redshift higher than 0.9. As a consistent test, we compared with the DES Collaboration, because our and their results diverge slightly for higher redshift, we conclude that this algorithm is sensitive to the training set available especially in the high redshift end.
The other test we conducted involved selecting the best galaxies according to their redshift PDFs. The resulting sub-samples gave BPZ with the biggest number of selected objects. However, because of the Bayesian method, this estimator forces smoother PDFs. ANNz2 shows better precision despite not having a sample as big as BPZ’s. ENF has the worst selection of Gaussian PDFs, with very few galaxies left for an LSS study. Lastly, worst in precision lies DNF but with sufficient galaxies for cosmological analysis. For the Small Peaks sub-sample, ANNz2 performed better than the other algorithms. Due to the highly selective criteria, the Gaussian sub-sample is not ideal for studies that need significant statistics, this will be demonstrated in a following study using Pz Cats.
@@
Lastly, we analysed how the bias behaves with faint galaxies. The result confirmed the other analysis, where ANNz2 is consistent for all redshifts, while the other estimators overestimate or underestimate the photo-zs. In the end, of the four algorithms ANNz2 was successful in most of the metrics discussed with less catastrophic results. In the next study, we will show whether this success influences cosmological results.

### v3

@@ 4.2 Precision with respect to the redshift
Complementary to the results above, we show $\sigma_{68}$ w.r.t. the spectroscopic redshift of the matched galaxies used for testing (all galaxies for BPZ). The result is shown in Figure 13. BPZ in figure 13(a) looses precision when $z>1.0$, for ENF/DNF(figs. 13(c),13(d)) this happens at $z>0.9$. The only photo-z estimator to show consistent and precise results is ANNz2.
@@
Figure 13: $\sigma^{68}$ vs $z_{spec}$ bin. Blue: positive error. Red: negative error.
@@
Table 4: Catastrophic results w.r.t the PDF mode for each algorithm.
We compared the relative frequency of galaxies classified with Gaussian PDFs from the testing sample for each algorithm. In Figure 14, we show the number of Gaussians ($N_{Gauss}$) up to $|\sigma_{z}|$ compared to the number of galaxies in the test sample up to $|\sigma_{z}|$. BPZ has nearly 70% Gaussian PDFs with bias bigger than 0.075 and ANNz2 has nearly 60%. ENF reaches a maximum of 25% of Gaussian PDFs for very small bias, after $|\sigma_{z}|>0.025$, the frequency saturates and only 15% of the PDFs are nearly Gaussian. The opposite happens to DNF, it is an improvement compared to ENF, but saturation is below 25%. This result shows an important situation that may occur; even though the PDFs are smooth, there are catastrophic redshift results.
Figure 14: Relative frequency of Gaussian distributions compared to the total number of testing samples w.r.t. the absolute bias ($\sigma_{z}$) for each estimator. BPZ: black, ANNz2: blue, ENF: gray, and DNF: red.
@@ 4.3 Filters with respect to the bias
Figure 15: Redshift ias distribution for each filter normalized by $\sigma^{68}$. Red: DNF, black: BPZ, gray: ENF and blue: ANNz2.
@@ 4.4 Colour cut
In the previous test, we based our $\sigma_{68}$ analysis w.r.t. the colour coverage of our spectroscopic sample, in Figure 1. For each estimator, we cut the redshifts restriction to the colours $41<g-i<$ 4 and $1.40.5<r-i<$ 1.4. Then, we selected the redshifts whose $\sigma_{68}$ is smaller than $10\%$ of the photo-z value. In table 5, there are the $z_{\rm eff}$ considering the colour cut and it is clear that they are compatible with the entire sample photo-z estimation. lastly, the number of objects reduces significantly to DNF, so depending on the use of the catalog, one must understand that an accurate representation of the filters by the training sample may reduce statistics.
Estimator $z_{\rm eff}$ number of galaxies
@@
We checked the bias distribution normalised by $\sigma_{68}$ in terms of the faintest part of each filter. This is shown in Figure 15, the only algorithm that is symmetric on the bias is ANNz2, BPZ underestimates the faint galaxies, while DNF and ENF overestimate them.
Fortunately, the criteria we have established for probability density functions have shown to be highly effective in eliminating suboptimal spectroscopic sample representations based on colour. This is evidenced in Figure 9 where we plot the $68\%$ confidence level of color surface distribution, which shows a notable reduction in subrepresented samples once we exclude the PDFs associated with high noise levels. The decrease in such subrepresentation becomes even more pronounced when we apply the Gaussian cut criterion, as depicted in Fig. 10. Within this Gaussian framework, the algorithm ANNz2 emerges as the most accurate among the methods evaluated, whereas ENF experiences a substantial reduction, although this comes at the cost of some statistical accuracy. Ultimately, our careful selection strategy guarantees a reliable representation with respect to colour metrics, ensuring the overall robustness of our analytical approach.
[…]
@@ 5 Summary
We selected 25,760 galaxies from four different spectroscopic surveys and cross-matched them with the DES Y3 BAO sample. These galaxies served to understand the redshift bias $\sigma$ and its 68th percentile $\sigma_{68}$ and are representative in terms of magnitude. We found that within a range of $0.850.79<z_{p}<$ 0.85 there is the lowest $\sigma$ for all the estimators we analysed. DNF has the biggest absolute value of the bias, while ENF, ANNz2 and BPZ lose precision for a redshift range below 0.7 and greater than 0.9.
@@
We also investigated the $\sigma_{68}$ metric, a robust metric to estimate the error of the algorithms. ANNz2 and BPZ showed consistent $\sigma_{68}$ for all bins, while ENF and DNF have an asymmetric error distribution, with lower negative error values than positive ones and a considerable error increase for redshift greater than 0.9. As a consistent test, we compared with the DES Collaboration, because our and their results diverge slightly for higher redshift, we conclude that this algorithm is sensitive to the training set available especially in the high redshift end.
In another evaluation we conducted, our objective was to identify the optimal galaxies based on their redshift Probability Density Functions (PDFs). This endeavour resulted in a subset where BPZ identified the highest count of objects. Nevertheless, the Bayesian methodology intrinsic to BPZ tends to produce more homogenised PDFs. Conversely, ANNz2, although possessing a smaller subset than BPZ, demonstrated superior accuracy. On the other hand, ENF generated the least effective selection, associated with Gaussian PDFs, leaving a minimal quantity of galaxies suitable for a Large-Scale Structure (LSS) analysis, only 14 galaxies per degree squared. DNF, while being the least precise of the methods, still retained an adequate number of galaxies for cosmological investigations. Within the Small Peaks sub-sample, ANNz2 outshone the other systems. Given the stringent criteria of the Gaussian subsample, it is less favourable for analyses requiring robust statistical outcomes. A noticeable effect of selecting PDFs is an improved representation in terms of colours compared with the spectroscopic sample; this finding underscores that such selection is significantly influenced by the magnitude space of the sample.
@@
Finally, we conducted an analysis of how bias manifests in the context of faint galaxies. Our findings corroborate the previous analyses, indicating that ANNz2 maintains consistency across all tested redshift ranges, whereas the other estimators exhibit tendencies to either overestimate or underestimate the photometric redshifts. Ultimately, among the four algorithms evaluated, ANNz2 demonstrated success across the majority of the metrics we discussed, with comparatively fewer catastrophic failures. In an upcoming study, we aim to investigate whether this level of success exerts an influence on cosmological outcomes.
6 Code availability
The particular software packages used in this work will be accessible at https://github.com/psilvaf/bao_pz and https://github.com/psilvaf/cat_org.
@@ Acknowledgments
The authors would like to thank the anonymous referee who provided useful and detailed comments.